OpenStudy (tabithax3):

how do you find the domain of (x-3)^-1/4?

OpenStudy (campbell_st):

well you are rally looking at \[\frac{1}{\sqrt[4]{x - 3}}\] so x cannot equal 3 as this would result in a zero denominator... and the have a real number in the denominator x > 3 would be the domain.

OpenStudy (tabithax3):

@campbell_st the problem is \[h(x)=\left( x-3 \right)^{-1/4}\] so it means\[-\frac{ 1 }{ \sqrt[4]{x-3} }\] right? and the domain is \[\left( 3,\infty \right)\] or do i ignore the - in \[(x-3)^{-1/4}\]

OpenStudy (tabithax3):

i hope that made sense...

OpenStudy (campbell_st):

if you have \[h(x) = (x - 3)^\frac{-1}{4}\] then it can be written as \[h(x) = \frac{1}{\sqrt[4]{x - 3}}\] so if you are dealing with real numbers then x cannot be less than 3 as it result in the 4th root of a negative... which means complex numbers... so you need to look at the set of numbers where x > 3 or \[(3, \infty)\] this set of numbers will result in a real number answer. hope this helps

OpenStudy (tabithax3):

got it, thank you for the help!